# appAppendix References (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v1

@@ Abstract
The diffusion model has long been plagued by scalability and quadratic complexity issues, especially within transformer-based structures. In this study, we aim to leverage the long sequence modeling capability of a State-Space Model called Mamba to extend its applicability to visual data generation. Firstly, we identify a critical oversight in most current Mamba-based vision methods, namely the lack of consideration for spatial continuity in the scan scheme of Mamba. Secondly, building upon this insight, we introduce a simple, plug-and-play, zero-parameter method named Zigzag Mamba, which outperforms Mamba-based baselines and demonstrates improved speed and memory utilization compared to transformer-based baselines. Lastly, we integrate Zigzag Mamba with the Stochastic Interpolant framework to investigate the scalability of the model on large-resolution visual datasets, such as FacesHQ $1024\times 1024$ and UCF101, MultiModal-CelebA-HQ, and MS COCO $256\times 256$. Code will be released at https://taohu.me/zigma/
@@ 1 Introduction
Diffusion models have demonstrated significant advancements across various applications, including image processing [68], video analysis [39], point cloud processing [88], and human pose estimation [28]. Many of these models are built upon Latent Diffusion Models (LDM)[68], which are typically based on the UNet backbone. However, scalability remains a significant challenge in LDMs[44]. Recently, transformer-based structures have gained popularity due to their scalability [66, 8] and effectiveness in multi-modal training [9]. Notably, the transformer-based structure DiT [66] has even contributed to enhancing the high-fidelity video generation model SORA [65] by OpenAI. Despite efforts to alleviate the quadratic complexity of the attention mechanism through techniques such as windowing [60], sliding [11], sparsification [16, 49], and hashing [17, 75], it remains a bottleneck for diffusion models.
@@ 2 Related Works
SDE and ODE in Diffusion models. The realm of Score-based Generative Models encompasses significant contributions from foundational works such as Score Matching with Langevin Dynamics (SMLD) by Song et al. [72] and the advent of Denoising Diffusion Probabilistic Models(DDPM) [38, 71]. These methodologies operate within the framework of Stochastic Differential Equations (SDEs), a concept further refined in the research of Song et al. [73]. Recent research strides, as exemplified by Karras et al. [45] and Lee et al. [50], have showcased the efficacy of employing Ordinary Differential Equation (ODE) samplers for diffusion SDEs, offering significant reductions in sampling costs compared to traditional approaches that entail discretizing diffusion SDEs. Furthermore, within the domain of Flow Matching methodologies [54, 57, 3, 4], both SMLD and DDPMs emerge as specialized instances under distinct paths of the Probability Flow ODE framework [73]. These models typically utilize velocity field parameterizations employing the linear interpolant, a concept that finds broader applications in the Stochastic Interpolant framework [3], with subsequent generalizations extending to manifold settings [12]. Explored in various topics such as image generation [54, 40], editing [43], video generation [20], motion generation [42], point cloud generation [88], Reinforcement Learning [96] and text generation [41]. The SiT model [63] scrutinizes the interplay between interpolation methods in both sampling and training contexts, albeit in the context of smaller resolutions such as $512\times 512$. Our research endeavors to extend these insights to a larger scale, focusing on the generalization capabilities for 2D images of $1024\times 1024$ and 3D video data.
@@ 3.2 Diffusion Backbone: Zigzag Mamba
(c) Factorized 3D Zigzag = 2D Zigzag + 1D Sweep: To address the suboptimal optimization issue, we propose to factorize the spatial and temporal correlations as separate Mamba blocks. The order of their application can be adjusted as desired, for example, "sstt" or "ststst", where "s" represents the spatial-zigzag Mamba and "t" represents the temporal-zigzag Mamba.
@@ 4.2 Ablation Study
Order Receptive Field. We propose a new concept in Mamba-based structure for non-1D data. Given that various spatially-continuous zigzag paths may exist in non-1D data, we introduce the term Order Receptive Field which denotes the number of zigzag paths explicitly employed in the network design.
@@ 5 Conclusion
In this paper, we present the Zigzag Mamba Diffusion Model, developed within the Stochastic Interpolant framework. Our initial focus is on addressing the critical issue of spatial continuity. We then devise a Zigzag Mamba block to better utilize the inductive bias in 2D images. Further, we factorize the 3D Mamba into 2D and 1D Zigzag Mamba to facilitate optimization. We empirically design various ablation studies to examine different factors. This approach allows for a more in-depth exploration of the Stochastic Interpolant theory. We hope our endeavor can inspire further exploration in the Mamba network design. In the future, we plan to examine this network design for discriminative tasks.

### v2

@@ Abstract
The diffusion model has long been plagued by scalability and quadratic complexity issues, especially within transformer-based structures. In this study, we aim to leverage the long sequence modeling capability of a State-Space Model called Mamba to extend its applicability to visual data generation. Firstly, we identify a critical oversight in most current Mamba-based vision methods, namely the lack of consideration for spatial continuity in the scan scheme of Mamba. Secondly, building upon this insight, we introduce Zigzag Mamba, a simple, plug-and-play, minimal-parameter burden, DiT style solution, which outperforms Mamba-based baselines and demonstrates improved speed and memory utilization compared to transformer-based baselines. Lastly, we integrate Zigzag Mamba with the Stochastic Interpolant framework to investigate the scalability of the model on large-resolution visual datasets, such as FacesHQ $1024\times 1024$ and UCF101, MultiModal-CelebA-HQ, and MS COCO $256\times 256$.
@@ 1 Introduction
Diffusion models have demonstrated significant advancements across various applications, including image processing [67], video analysis [40], point cloud processing [87], and human pose estimation [29]. Many of these models are built upon Latent Diffusion Models (LDM)[67], which are typically based on the UNet backbone. However, scalability remains a significant challenge in LDMs[41]. Recently, transformer-based structures have gained popularity due to their scalability [65, 9] and effectiveness in multi-modal training [10]. Notably, the transformer-based structure DiT [65] has even contributed to enhancing the high-fidelity video generation model SORA [64] by OpenAI. Despite efforts to alleviate the quadratic complexity of the attention mechanism through techniques such as windowing [58], sliding [12], sparsification [18, 46], hashing [19, 74], Ring Attention [53, 14], Flash Attention [21] or a combination of them [8, 97], it remains a bottleneck for diffusion models.
@@ 2 Related Works
SDE and ODE in Diffusion models. The realm of Score-based Generative Models encompasses significant contributions from foundational works such as Score Matching with Langevin Dynamics (SMLD) by Song et al.[71], and the advent of Diffusion Models with Denoising Score Matching (DDPMs) proposed by Ho et al.[39]. These methodologies operate within the framework of Stochastic Differential Equations (SDEs), a concept further refined in the research of Song et al. [72]. Recent research strides, as exemplified by Karras et al.[42] and Lee et al.[47], have showcased the efficacy of employing Ordinary Differential Equation (ODE) samplers for diffusion SDEs, offering significant reductions in sampling costs compared to traditional approaches that entail discretizing diffusion SDEs. Furthermore, within the domain of Flow Matching methodologies [51] and Rectified Flow architectures [55], both SMLD and DDPMs emerge as specialized instances under distinct paths of the Probability Flow ODE framework [72]. These models typically utilize velocity field parameterizations employing the linear interpolant, a concept that finds broader applications in the Stochastic Interpolant framework [3], with subsequent generalizations extending to manifold settings [13]. The SiT model [61] scrutinizes the interplay between interpolation methods in both sampling and training contexts, albeit in the context of smaller resolutions such as $512\times 512$. Our research endeavors to extend these insights to a larger scale, focusing on the generalization capabilities for 2D images of $1024\times 1024$ and 3D video data.
@@ 3.2 Diffusion Backbone: Zigzag Mamba
(c) Factorized 3D Zigzag = 2D Zigzag + 1D Sweep: To address the suboptimal optimization issue, we propose to factorize the spatial and temporal correlations as separate Mamba blocks. The order of their application can be adjusted as desired, for example, "sstt" or "ststst", where "s" represents the spatial-zigzag Mamba and "t" represents the temporal-zigzag Mamba. For a 1D temporal sweep, we simply opt for forward and backward scanning, since there is only one dimension on the time axis.
@@ 4.2 Ablation Study
Order Receptive Field. We propose a new concept in Mamba-based structure for multidimensional data. Given that various spatially-continuous zigzag paths may exist in multidimensional data, we introduce the term Order Receptive Field which denotes the number of zigzag paths explicitly employed in the network design.
@@ 5 Conclusion
In this paper, we present the Zigzag Mamba Diffusion Model, developed within the Stochastic Interpolant framework. Our initial focus is on addressing the critical issue of spatial continuity. We then devise a Zigzag Mamba block to better utilize the inductive bias in 2D images. Further, we factorize the 3D Mamba into 2D and 1D Zigzag Mamba to facilitate optimization. We empirically design various ablation studies to examine different factors. This approach allows for a more in-depth exploration of the Stochastic Interpolant theory. We hope our endeavor can inspire further exploration in the Mamba network design.